One number is 1 less than another. If 6 times the smallest number minus 2 times the larger number is 30, find the two numbers.
Thanks:)

Answer by scott8148(5891)

let s=smallest, so s+1=largest ... 6(s)-2(s+1)=30 ... 6s-2s-2=30 ... 4s=32 ... s=8 ... s+1=9